OpenStudy (anonymous):

Can someone explain to me how the indefinite integral of squareroot of x over x is 2 squareroot of x?

OpenStudy (anonymous):

This is my function \[\int\limits_{}^{}\frac{ 1+\sqrt{x} }{ x }dx\]

OpenStudy (anonymous):

I get how 1/x is lnx but I'm not sure how wolfram alpha is getting, \[2\]

OpenStudy (anonymous):

\[2\sqrt{x}\]

OpenStudy (inkyvoyd):

What did you type in wolfram alpha?

OpenStudy (anonymous):

integrate (1+x^(1/2))/(x)

OpenStudy (inkyvoyd):

you're integrating sqrt(x)/x=1/sqrt(x) for that second part

OpenStudy (inkyvoyd):

1/sqrt(x)=x^(-1/2) divide by -1/2+1 and take x to the power of -1/2+1 according to power rule and you get 2x^(1/2)=2sqrt(x)

OpenStudy (inkyvoyd):

@JerJason ?

OpenStudy (anonymous):

sorry i had to leave for a moment.

OpenStudy (anonymous):

How is sqrt(x)/x=1/sqrt(x)?

OpenStudy (loser66):